OpenStudy (anonymous):

Find cos θ if sin θ =-12/13 and tan θ > 0

OpenStudy (jdoe0001):

\(\bf sin(\theta) = -\cfrac{12}{13} \implies \cfrac{\textit{opposite side}}{\textit{hypotenuse}} \implies \cfrac{b}{c}\\ \textit{keep in mind that}\\ c^2 = a^2+b^2 \implies \sqrt{c^2-b^2} = a^2\\ \textit{if } tan(\theta) > 0\\ tan(\theta) = \cfrac{\textit{opposite side}}{\textit{adjacent side}}\implies\cfrac{b}{a}\\ sin(\theta) \textit{ is negative, thus }cos(\theta) \textit{ is also negative}\)

OpenStudy (dumbcow):

there are 2 ways to go about this... i will show 1 way which is using formula \[\sin^{2}x + \cos^{2}x = 1\] ...oh nevermind

OpenStudy (anonymous):

Okay, so then you just plug in the sides?

OpenStudy (campbell_st):

its a 3rd quadrant angle, tan > 0 sin <0 draw a picture  find x using pythagoras' theorem and it will be negative because of its location on the number plane. then you can find cos